OpenStudy (anonymous):

find all the eighth roots of (19+7i)

OpenStudy (perl):

solve z^8= 19+ 7i , where z is a complex number

OpenStudy (fibonaccichick666):

Demouvre?

OpenStudy (perl):

yes

OpenStudy (fibonaccichick666):

(that is a hint as well as a question on my part :) thanks @perl)

OpenStudy (perl):

This is a very useful theorem , which we can use here.

OpenStudy (anonymous):

did you know answer of this question

OpenStudy (anonymous):

yes

OpenStudy (perl):

lets start with Demoivre's theorem

OpenStudy (fibonaccichick666):

please work with perl @vikiy , it is against the code of conduct to just want solutions. We are all here to help, not just give answers

OpenStudy (perl):

and then you can solve it :)

OpenStudy (perl):

if anybody wants to post Demoivre's theorem, please don't hesitate

ganeshie8 (ganeshie8):

19+7i is a messy number

OpenStudy (perl):

roughly speaking ( r cos t + i sin t ) ^ (n) = r^n ( cos nt + i sin nt)

OpenStudy (perl):

so the trick is to get 19 + 7i into r cos t + i sin t , form

OpenStudy (perl):

r = sqrt( 19^ + 7^2) t = atan( 19,7)